Use the exponential decay model A=Pe^(kt) to solve the following proublem.
The half-life of plutonium-238 is days. The initial amount is 100 grams.
a)find the value of K to four places.
b) write the formula for the amount of plutonium-238 present at the end of T days.
c)How much of the initial amount will remain after 30 days? Round the answer to the nearest tenth gram.
I know this seems like alot but I swear it's all one proublem just has Three parts to it, been having trouble, thank you sooo much in advance!!! :)

Answer by stanbon(57246)

Use the exponential decay model A=Pe^(kt) to solve the following proublem.
The half-life of plutonium-238 is days. The initial amount is 100 grams.
a)find the value of K to four places.
(1/2)P = P*e^(k*238)
e^(238k) = 1/2
Take the natural log of both sides to get:
238k = ln(1/2)
k = -0.002912
---------------------------
b) write the formula for the amount of plutonium-238 present at the end of T days.
A(t) = P*e^(-0.002912t)
---------------------------
c)How much of the initial amount will remain after 30 days? Round the answer to the nearest tenth gram.
A(30) = 100*e^(-0.002912*30
---
A(30) = 100*0.916337
A(30) = 91.6337 grams
==========================
Cheers,
Stan H.
=============